As an economics student your budget for meals at the university is Rs 12,000 and you have two options during the course of you studies:


 


Eating at the university cafeteria for Rs 40 per meal

Eating at the dining hall for Rs 10 per meal.

 


a. Draw the budget constraint showing the trade-off between cafeteria and dinning hall meals, with cafeteria meals on x-axis and dining hall meals on y-axis. Assume you spend equal amounts on both goods, draw an indifference curve showing the optimum choice. Label the optimum as point A.


1
Expert's answer
2021-06-01T11:28:41-0400

The vertical axis measures the quantity of dining hall meals. The horizontal axis measures the quantity of cafeteria meals. The budget constraint line measures the combination of two meals the individual could afford to purchase. When the individual wish to consume 300 cafeteria meals during the courses, the individual could no longer consume dining hall meals. When the individual wish to consume 1200 dining hall meals during the courses, the individual could no longer consume cafeteria meals. The individual could purchase the combinations of two meals that are traced by the budget constraints line. The indifference curve suggests that the satisfaction in consuming a combination of two goods remains the same in all the points of the indifference curve. 

There is only one optimal point i.e., point A where the budget constraint line is tangent to the indifference curve. This point indicates that the marginal rate of substitution is equal to the opportunity cost of two goods.

"quantity \space received= \frac {income}{ cafeteria \space price
\space per\space meal}"

"=\frac{12000}{40}=300"

"quantity \space received= \frac {income}{ dinning \space price
\space per\space meal}"

"=\frac{12000}{10}=1200"